Alan Madison needs $244,400 in 10 years.

Click here to view factor tables

How much must he invest at the end of each year, at 5% interest, to meet his needs? (Round factor values to 5 decimal places, e.g. 1.25124 and final answers to 0 decimal places, e.g. 458,581.)
Answer #1
Please give positive ratings so I can keep answering. It would help me a lot. Please comment if you have any query. Thanks!
Alan Madison
Future Value of Annuity P*((1+i)^n-1)/i
Here P= ?
Future Value of Annuity= 244,400.00
i= 5%=0.05
n=             10.00
244400= P*((1+0.05)^10-1)/.05
244400= P*((1.05)^10-1)/.05
244400= P*(1.62889-1)/.05
244400= P*(0.62889)/.05
244400= P*12.5778
P= 244400/12.5778
P=     19,431.06
So Alan must invest $ 19,431 at the end of each year.